how do i figure out this simple problem?
25 divided by .5?

Answer by richwmiller(17219)

There are many ways
1-we could use a calculator
2- 25/.5= 250/5=50
We could change .5 to 1/2
3-25/1/2=25*2=50
Remember multiplying by the reciprocal is the same as dividing by the number.
The reciprocal of 1/2 is 2
So dividing by 1/2 is the same as multiplying by 2